# Rethinking 3D Dense Caption and Visual Grounding in A Unified Framework through Prompt-based Localization

### v1

@@ Abstract
3D Visual Grounding (3DVG) and 3D Dense Captioning (3DDC) are two crucial tasks in various 3D applications, which require both shared and complementary information in localization and visual-language relationships. Therefore, existing approaches adopt the two-stage “detect-then-describe/discriminate” pipeline, which relies heavily on the performance of the detector, resulting in suboptimal performance. Inspired by DETR, we propose a unified framework, 3DGCTR, to jointly solve these two distinct but closely related tasks in an end-to-end fashion. The key idea is to reconsider the prompt-based localization ability of the 3DVG model. In this way, the 3DVG model with a well-designed prompt as input can assist the 3DDC task by extracting localization information from the prompt. In terms of implementation, we integrate a Lightweight Caption Head into the existing 3DVG network with a Caption Text Prompt as a connection, effectively harnessing the existing 3DVG model’s inherent localization capacity, thereby boosting 3DDC capability. This integration facilitates simultaneous multi-task training on both tasks, mutually enhancing their performance. Extensive experimental results demonstrate the effectiveness of this approach. Specifically, on the ScanRefer dataset, 3DGCTR surpasses the state-of-the-art 3DDC method by 4.30% in CIDEr@0.5IoU in MLE training and improves upon the SOTA 3DVG method by 3.16% in Acc@0.25IoU.
Figure 1: Illustration of existing method (upper) with two-stage pipeline and our single-stage 3DGCTR (bottom). These two-stage methods heavily depend on a detector’s output and also suffer from low reuse of task-agnostic modules. Therefore, we propose a transformer-based single-stage model that simply builds upon a mature 3DVG model, thus giving 3DVG model 3DDC capability. Compared to the SOTA method 3DJCG (Cai et al., 2022) that jointly trains the two tasks, our method achieves a significant improvement.
@@ 1 Introduction
The intersection of 3D scene understanding and natural language processing has emerged as a focal point in research, evident in tasks like 3D visual grounding (3DVG) (Guo et al., 2023; Zhang et al., 2023; Hsu et al., 2023; Chen et al., 2022; Jain et al., 2022; Wu et al., 2023; Zhao et al., 2021; Yang et al., 2021; Yuan et al., 2021; Roh et al., 2022; Luo et al., 2022; Yin et al., 2023; Huang et al., 2021) and 3D dense captioning (3DDC) (Yuan et al., 2022; Chen et al., 2021b; Jiao et al., 2022; Zhong et al., 2022a; Yu et al., 2023). 3DVG involves processing a text paired with a point cloud to identify the specified object with a bounding box. Its counterpart, 3DDC, takes a point cloud and produces detailed bounding boxes and descriptions for each object. These tasks are pivotal for developing innovative applications, including assistive robotics and intuitive language-based interaction in AR/VR environments.
@@
Considering that 3DVG and 3DDC contain both shared and complementary information in nature, previous works (Chen et al., 2021a; Cai et al., 2022; Chen et al., 2023b) attempt to integrate these two tasks into a unified work. D3Net (Chen et al., 2021a) proposes a speaker-listener pipeline to improve the performance of these two tasks in a self-critical manner. Based on a pre-trained detector, 3DJCG (Cai et al., 2022) utilizes an attribute and relation-aware module to enhance the task-agnostic features. Unit3D (Chen et al., 2023b) propose a transformer-based fusion module that is based on a PointGroup (Jiang et al., 2020) detection backbone and a BERT (Devlin et al., 2018) encoder, to learn the multi-model representations between objects and text inputs. Among existing methods, they all adopt a detector-based architecture (Chen et al., 2021a; Cai et al., 2022; Chen et al., 2023b), in which two task-specific modules were stacked on a shared detector (Ding et al., 2019; Jiang et al., 2020) to unify these two tasks in a single framework. Though these two-stage methods have achieved remarkable performance, the detector-based architecture yields suboptimal performance due to the following issues: 1) For 3DVG, the language-irrelevant proposals generated by the detector easily cause incorrect localization, which is demonstrated by 3D-SPS (Luo et al., 2022). 2) For 3DDC, the serial and explicit reasoning highly limits the mutual promotion of localization and captioning.
@@
Recently, DETR-like (Carion et al., 2020) models developed based on Transformers (Vaswani et al., 2017) have achieved inspiring progress on both 3DVG (Luo et al., 2022; Jain et al., 2022; Wu et al., 2023) and 3DDC (Chen et al., 2023a) tasks in a single-stage way. It is a natural idea to integrate these two tasks into a single DETR-like architecture, thus maximizing the use of task-agnostic modules to achieve end-to-end training. However, this idea is not easy to implement due to the following reasons: 1) Variations in Input Types. In terms of inputs, 3DDC exclusively requires a point cloud from a scene, whereas 3DVG also demands additional text references. 2) Divergent Optimization Objectives for Query Embeddings. Query embeddings are an important component of the DETR-like architecture, which represents the object proposals to be located in a scene. In 3DVG methods (Luo et al., 2022; Jain et al., 2022; Wu et al., 2023), the query embeddings are learned to align the target object described in the input referencing text. In contrast, 3DDC (Chen et al., 2023a) focuses on aligning query embeddings with all predetermined ground-truth objects in the scene.
@@
To address the above issues and naturally merge two different tasks, we rethink that a well-crafted grounding referencing text can act as a prompt to boost the detection capability of the 3DVG model, which is also supported by the previous work (Jain et al., 2022). By utilizing such a prompt as a link, it is possible to cast the detection part of the dense caption model as referential grounding. Specifically, as described above, the query embeddings are trained to align the input text in the 3DVG task. When a prompt such as “cabinet . bed . chair . sofa . ” composed of multiple object class names is fed into the 3DVG model, the 3DVG model can locate the objects mentioned in the prompt, thus turning it into a detector. This allows us to naturally merge the two tasks into the DETR-like architecture by harmonizing the input form across both tasks (point cloud together with prompt) and the optimization objectives for query embeddings (align the objects mentioned in the prompt). Combined with our rethinking, we proposed a unified Transformer framework termed 3DGCTR, which integrates the Dual-Clued Captioner (DCC), a lightweight caption head outlined in (Chen et al., 2023a), into the 3DVG model. Together with different text prompts for the two tasks, we enable the model to train both tasks end-to-end.
@@
Experiments show that both the 3DDC and 3DVG performance of 3DGCTR have achieved state-of-the-art on the ScanRefer(Chen and Chang, 2020) benchmark. To be specific, 3DGCTR surpasses the 3DDC method by 4.30% in CIDEr@0.5IoU in MLE training and improves upon the 3DVG method by 3.16% in Acc@0.25IoU. Meanwhile, through joint training, 3DGCTR can achieve the mutual promotion of the two tasks. Specifically, the 3DDC and 3DVG tasks increased CIDEr@0.5IoU by 1.27% and Acc@0.25IoU by 0.3%, respectively.
@@
  • Rethinking the role of prompts in the 3DVG model. We provide a new perspective on the prompt-based localization ability of the 3DVG model, transforming it into a dual-purpose tool that effectively facilitates both 3DVG and 3DDC tasks.
  • Advanced Unified Framework for 3D Visual Grounding and 3D Dense Captioning. By adding a lightweight caption head to the 3DVG model and using a well-designed prompt for the 3DDC task, our 3DGCTR framework integrates 3DVG and 3DDC tasks within a DETR-like architecture, enabling efficient end-to-end training.
@@
Figure 2: The framework of 3DGCTR builds upon a mature DETR-like 3DVG model (bottom in the figure) with a lightweight caption head. After obtaining the fused visual tokens $V$ and decoder output query embeddings $Q$ of each scene, the caption head uses $Q$ as caption prefix to identify the described region, and contextual features $V$ surrounding the vote query to complement with more surrounding information for more descriptive caption generation. Finally, the referring/detection boxes are selected from the candidate boxes via the referring scores.
@@ 2 Related Work
Two-Stage Method. Within the 3DREC realm, most recent works (Zhao et al., 2021; Yang et al., 2021) employ a two-stage pipeline. Initially, they utilize either ground truth or a 3D object detector (Jiang et al., 2020; Liu et al., 2021) for generating object proposals. Subsequently, these methods use text and 3D encoder (Liu et al., 2019; Qi et al., 2017) to extract features and then ground the target one after the feature fusion. For instance, InstanceRefer (Yuan et al., 2021) ingeniously construes the task as instance-matching, while LanguageRefer (Roh et al., 2022) reinterprets it as a linguistic modeling task by replacing the 3D features with predicted object labels.
Single-Stage Method. 3D-SPS (Luo et al., 2022) is the first to introduce a single-stage network for 3DREC. Meanwhile, recent state-of-the-art 3DREC approaches, such as BUTD-DETR (Jain et al., 2022) and EDA (Wu et al., 2023), have incorporated the query selection modules from Groupfree (Liu et al., 2021), indicating a converging trend of these technologies. It is worth noting that the two-stage setting introduced in (Wu et al., 2023) is different from that described in our paper. Specifically, the outputs of the detector participate in the cross-attention of the query embedding as auxiliary information, which in essence still belongs to the single-stage training method. In comparison, the 3DRES field remains largely unexplored, with the only notable single-stage method TGNN (Huang et al., 2021). TGNN builds on an instance segmentation framework, employing text-guided Graph Neural Networks to pinpoint the center of the instance referred to in the text. In this paper, we will focus our research on 3DREC because it has been fully explored.
@@
3D dense captioning is a task that requires translating 3D scene information to a set of bounding boxes and natural language descriptions, is challenging and has raised great interest among scholars in recent years.
Two-Stage Method. Scan2Cap (Chen et al., 2021b) and MORE (Jiao et al., 2022) create graphs based on a detector’s (Qi et al., 2019; Jiang et al., 2020) box predictions, using predefined rules to decipher complex object relations in 3D scenes. SpaCap3D (Wang et al., 2022a) develops a spatially-informed transformer to understand spatial relationships among detector outputs. 3DJCG (Cai et al., 2022) and D3Net (Chen et al., 2021a) explore the mutual enhancement of 3D dense captioning and 3D visual grounding. $\mathcal{X}$-Trans2Cap (Yuan et al., 2022) incorporates 2D priors to enhance 3D dense captioning through knowledge transfer. Recently, (Zhong et al., 2022b) has focused on contextual data for recognizing non-object details. These methods have significantly advanced the 3D dense captioning challenge. Yet, they are heavily reliant on the accuracy of the detector.
Single-Stage Method. Vote2Cap-DETR (Chen et al., 2023a) is a unified, single-stage approach that concurrently detects objects and generates captions, addressing 3D dense captioning as a set prediction issue. Mirroring Vote2Cap-DETR (Chen et al., 2023a), our 3DGCTR model seamlessly integrates 3DDC and 3DVG tasks within a DETR-like framework. Furthermore, 3DGCTR’s visual grounding aspect enhances the dense captioning task more effectively than joint training of detection and captioning tasks, owing to its superior text-visual alignment.
@@ 3 Method
The key idea of 3DGCTR is to harness the relation-oriented capability of the 3DVG task and the object-oriented ability of the 3DDC task to complement each other. The architecture of the adopted 3DVG model is illustrated in Sec. 3.1. To enhance the feature extraction ability, we replace the visual backbone in EDA (Wu et al., 2023) with PointMetaBased (Lin et al., 2023) as our 3DVG model, termed EDA-PMB, which is introduced in Sec. 3.2. To develop the 3DDC capacity, a lightweight caption head is appended to the 3DVG model for caption generation, which is detailed in Sec. 3.3. To naturally unify 3DDC and 3DVG tasks in DETR-like architecture, we propose a text prompt for 3DDC as an input to the 3DVG backbone, as described in Sec. 3.4. Last, the training scheme and the inference process are illustrated in Sec. 3.5 and Sec. 3.6, respectively.
@@ 3.1 Preliminaries: 3DVG Model
To ensure the end-to-end training characteristics of our model, we introduce EDA (Wu et al., 2023) with the setting of single-stage as the basic 3DVG model. EDA is a typical DETR-like (Carion et al., 2020) model composed of a backbone, a cross-encoder, and an object decoder.
Visual and Textual Backbone. As shown in Figure. 2, PointMetaBased (Lin et al., 2023) produces the visual tokens $V_{0}\in\mathbb{R}^{n\times d}$ from the input point cloud $P\in\mathbb{R}^{N\times 3}$, while Roberta-base (Liu et al., 2019) processes referential text for text tokens $T_{0}\in\mathbb{R}^{l\times d}$, where $n$ denotes the number of visual token, $l$ denotes the number of text token, $N$ denotes the point cloud size, and $d$ is the feature dimension.
Cross-Encoder. The visual tokens $V_{0}$ and text tokens $T_{0}$ enter a dual-pathway cross-encoder alternatively stacking with self-attention, cross-attention, and FFN layers, producing thoroughly fused features $V\in\mathbb{R}^{n\times d}$ and $T\in\mathbb{R}^{l\times d}$. The output visual tokens $V\in\mathbb{R}^{n\times d}$ is input to the KPS module proposed by (Liu et al., 2021). The top-k object queries $Q_{0}\in\mathbb{R}^{k\times d}$ are chosen and then fed into the object decoder, caption head and the query scoring branch.
@@ 3.2 Integrating PointMetaBase into EDA
A more powerful visual backbone can better extract the attributes and relative position relationships of objects in 3D scenes, which is helpful for more fine-grained interaction with language. Hence, we replaced the PointNet++ (Qi et al., 2017) in EDA (Wu et al., 2023) with our proposed PointMetaBase (Lin et al., 2023). The main contribution of PointMetaBase is to abstract the point cloud feature extraction network into a Meta Architecture consisting of four key modules and explore the optimal design of each module. However, as the K-Nearest query is used in the module Neighbor Update, the number of downsampled point clouds output by the backbone is not fixed, and thus can’t be used as the visual tokens $V_{0}$ directly. To solve this problem, we design a parameter-free visual token query module. We first use the farthest point sampling (FPS) algorithm to sample a fixed number of $n$ points from the original point cloud $P$ as the candidates $C$:
@@
Then, given the output tokens $V^{\prime}\in\mathbb{R}^{n^{\prime}\times 3}$ from the PointMetaBase, we employ the ball query (Qi et al., 2017) with a ball radius of $r$ to find $k_{q}$ nearest neighbor tokens in the coordinate space of $V^{\prime}$ for each candidate, then assign the features of $V^{\prime}$ to them using max-pooling. Finally, we can obtain the visual tokens $V_{0}\in\mathbb{R}^{n\times 3}$:
@@
3.3 Lightweight Caption Head
We argue that the key to unambiguous detailed caption generation is to obtain the relationship between the target object and its close surroundings. As described in Vote2Cap-DETR (Chen et al., 2023a), the vote queries fail to provide adequate attribute and spatial relations. To address this issue, they proposed a caption head named Dual-Clued Captioner (DCC), which introduced the vote queries’ $k$ nearest local context token features as their local surroundings and keys for cross attention. In our method, the top-k object queries $Q_{0}$ generated in the KPS module also have the same problem as the vote queries generated in Vote2Cap-DETR. However, different from Vote2Cap-DETR, where vote queries are constructed based on spatial bias and content information for better caption generation, our KPS-based method focuses on the precise localization of key points of objects to tackle both 3DVG and 3DDC tasks simultaneously. Therefore, we use DCC as our caption head with some modifications to fit this difference. Specifically, instead of using vote queries’ $k$ nearest for cross attention, we introduce the visual tokens $V$ that integrate the broader scene context as the object’s interaction and relation information with its surroundings, leading to richer and more contextually nuanced captions.
As shown in Figure 2, DCC is a lightweight transformer decoder-based model comprising two identical transformer decoder blocks, sinusoidal position embeddings $PE(\cdot)$, and a linear classification head. For effective caption generation, DCC processes dual visual cues $V_{c}=(Q,V)$. Firstly, following (Wang et al., 2022b), in captioning a proposal, the standard ’Start Of Sequences’ (SOS) prefix is replaced with the query $Q$ from the described query, identifying the object in focus. Then, for each query embedding $q\in Q$, DDC gives its corresponding dense caption. The masked self-attention and cross-attention mechanism of the transformer decodes the contextual relationship between the query and the visual scene, which can be formulated as:

### v2

@@ Abstract
3D Visual Grounding (3DVG) and 3D Dense Captioning (3DDC) are two crucial tasks in various 3D applications, which require both shared and complementary information in localization and visual-language relationships. Therefore, existing approaches adopt the two-stage “detect-then-describe/discriminate" pipeline, which relies heavily on the performance of the detector, resulting in suboptimal performance. Inspired by DETR, we propose a unified framework, 3DGCTR, to jointly solve these two distinct but closely related tasks in an end-to-end fashion. The key idea is to reconsider the prompt-localization ability of the 3DVG model. In this way, the 3DVG model with a well-designed prompt as input can assist the 3DDC task by extracting localization information from the prompt. In terms of implementation, we integrate a caption head into the existing 3DVG network with a caption text prompt as a connection, effectively harnessing the existing 3DVG model’s inherent localization capacity, thereby boosting 3DDC capability. This integration facilitates simultaneous multi-task training on both tasks, mutually enhancing their performance. Extensive experimental results demonstrate the effectiveness of this approach. Specifically, on the ScanRefer dataset, 3DGCTR surpasses the state-of-the-art 3DDC method by 4.30% in CIDEr@0.5IoU in MLE training and improves upon the SOTA 3DVG method by 3.16% in Acc@0.25IoU. The codes are at https://github.com/Leon1207/3DGCTR.
@@ 1 Introduction
The intersection of 3D scene understanding and natural language processing has emerged as a focal point in research, evident in tasks like 3D visual grounding (3DVG) [14, 40, 15, 7, 17, 34, 41, 35, 38, 29, 24, 36, 16] and 3D dense captioning (3DDC) [39, 9, 19, 42, 37]. 3DVG involves processing a text paired with a point cloud to identify the specified object with a bounding box. Its counterpart, 3DDC, takes a point cloud and produces detailed bounding boxes and descriptions for each object. These tasks are pivotal for developing innovative applications, including assistive robotics and intuitive language-based interaction in AR/VR environments.
@@
Considering that 3DVG and 3DDC contain both shared and complementary information in nature, previous works [6, 3, 10] attempt to integrate these two tasks into a unified work. D3Net [6] proposes a speaker-listener pipeline to improve the performance of these two tasks in a self-critical manner. Based on a pre-trained detector, 3DJCG [3] utilizes an attribute and relation-aware module to enhance the task-agnostic features. Unit3D [10] propose a transformer-based fusion module that is based on a PointGroup [18] detection backbone and a BERT [12] encoder, to learn the multi-model representations between objects and text inputs. Among existing methods, they all adopt a detector-based architecture [6, 3, 10], in which two task-specific modules were stacked on a shared detector [13, 18] to unify these two tasks in a single framework. Though these two-stage methods have achieved remarkable performance, the detector-based architecture yields suboptimal performance due to the following issues: 1) For 3DVG, the language-irrelevant proposals generated by the detector easily cause incorrect localization, which is demonstrated by 3D-SPS [24]. 2) For 3DDC, the serial and explicit reasoning highly limits the mutual promotion of localization and captioning.
Figure 1: Illustration of existing method (a) with two-stage pipeline and our single-stage 3DGCTR (b). Existing methods heavily depend on a detector’s output and also suffer from low reuse of task-agnostic modules. Therefore, we propose a transformer-based model that simply builds upon a mature 3DVG model, thus giving 3DVG model 3DDC capability. Compared to the SOTA method 3DJCG [3] that jointly trains the two tasks, our method achieves a significant improvement.
@@
Recently, DETR-like [4] models developed based on Transformers [30] have achieved inspiring progress on both 3DVG [24, 17, 34] and 3DDC [8] tasks in a single-stage way. It is a natural idea to integrate these two tasks into a single DETR-like architecture, thus maximizing the use of task-agnostic modules to achieve end-to-end training. However, this idea is not easy to implement due to the following reasons: 1) Variations in Input Types. In terms of inputs, 3DDC exclusively requires a point cloud from a scene, whereas 3DVG also demands additional text references. 2) Divergent Optimization Objectives for Query Embeddings. Query embeddings are an important component of the DETR-like architecture, which represents the object proposals to be located in a scene. In 3DVG methods [24, 17, 34], the query embeddings are learned to align the target object described in the input text. In contrast, 3DDC [8] focuses on aligning query embeddings with all predetermined ground-truth objects in the scene.
@@
To address the above issues and naturally merge two different tasks, we rethink the role of a well-crafted grounding referencing text. As described in [17] and [34], they leverage prompt-based detection for data augmentation. However, by utilizing such a prompt as a link, it is possible to cast the detection part of the dense caption model as referential grounding. Specifically, as described above, the query embeddings are trained to align the input text in the 3DVG task. When a prompt such as “cabinet . bed . chair . sofa . " composed of multiple object names is fed into the 3DVG model, the 3DVG model can locate the objects mentioned in the prompt, thus turning it into a detector. By harmonizing the input form and the optimization objectives for query embeddings across both tasks, it’s natural to merge the two tasks into the DETR-like architecture. This eliminates the two-stage optimization paradigm, where two tasks share visual representations and are optimized simultaneously during training, reinforcing each other. With this rethinking, we proposed a unified Transformer framework termed 3DGCTR, which integrates the Dual-Clued Captioner, a lightweight caption head outlined in [8], into the 3DVG model. Together with different text prompts for the two tasks, we enable the model to train both tasks end-to-end.
@@
Experiments show that both the 3DDC and 3DVG performance of 3DGCTR have achieved state-of-the-art on the ScanRefer[5] benchmark. To be specific, 3DGCTR surpasses the 3DDC method by 4.30% in CIDEr@0.5IoU in MLE training and improves upon the 3DVG method by 3.16% in Acc@0.25IoU. Meanwhile, through joint training, 3DGCTR can achieve the mutual promotion of the two tasks. Specifically, the 3DDC and 3DVG tasks increased CIDEr@0.5IoU by 1.27% and Acc@0.25IoU by 0.3%, respectively.
@@
  • Rethinking the role of prompts in the 3DVG model. We provide a new perspective on the prompt-localization ability of the 3DVG model, transforming it into a dual-purpose tool that effectively facilitates both 3DVG and 3DDC tasks.
  • Advanced Unified Framework for 3D Visual Grounding and 3D Dense Captioning. By adding a caption head to the 3DVG model and using a well-designed prompt for the 3DDC task, our 3DGCTR framework integrates 3DVG and 3DDC tasks within a DETR-like architecture, enabling efficient end-to-end training.
@@ 2 Related Work
Two-Stage Method. Within the 3DREC realm, most recent works [41, 35] employ a two-stage pipeline. Initially, they utilize either ground truth or a 3D object detector [18, 23] for generating object proposals. Subsequently, these methods use text and 3D encoder [22, 26] to extract features and then ground the target one after the feature fusion. For instance, InstanceRefer [38] ingeniously construes the task as instance-matching, while LanguageRefer [29] reinterprets it as a linguistic modeling task by replacing the 3D features with predicted object labels.
Single-Stage Method. 3D-SPS [24] is the first to introduce a single-stage network for 3DREC. Meanwhile, recent state-of-the-art 3DREC approaches, such as BUTD-DETR [17] and EDA [34], have incorporated the query selection modules from Groupfree [23], indicating a converging trend of these technologies. It is worth noting that the two-stage setting introduced in [34] is different from that described in our paper. Specifically, the outputs of the detector participate in the cross-attention of the query embedding as auxiliary information, which in essence still belongs to the single-stage training method. In comparison, the 3DRES field remains largely unexplored, with the only notable single-stage method TGNN [16]. TGNN builds on an instance segmentation framework, employing text-guided Graph Neural Networks to pinpoint the center of the instance referred to in the text. In this paper, we will focus our research on 3DREC because it has been fully explored.
@@
3D dense captioning is a task that requires translating 3D scene information to a set of bounding boxes and natural language descriptions.
Two-Stage Method. Scan2Cap [9] and MORE [19] create graphs based on a detector’s [27, 18] box predictions, using predefined rules to decipher complex object relations in 3D scenes. SpaCap3D [33] develops a spatially-informed transformer to understand spatial relationships among detector outputs. 3DJCG [3] and D3Net [6] explore the mutual enhancement of 3D dense captioning and 3D visual grounding. $\mathcal{X}$-Trans2Cap [39] incorporates 2D priors to enhance 3D dense captioning through knowledge transfer. Recently, [43] has focused on contextual data for recognizing non-object details. These methods have significantly advanced the 3D dense captioning challenge. Yet, they are heavily reliant on the accuracy of the detector.
Single-Stage Method. Vote2Cap-DETR [8] is a unified, single-stage approach that concurrently detects objects and generates captions, addressing 3D dense captioning as a set prediction issue. Mirroring Vote2Cap-DETR [8], our 3DGCTR model seamlessly integrates 3DDC and 3DVG tasks within a DETR-like framework. Furthermore, 3DGCTR’s visual grounding aspect enhances the dense captioning task more effectively than joint training of detection and captioning tasks, owing to its superior text-visual alignment.
Figure 2: The framework of 3DGCTR builds upon a mature DETR-like 3DVG model (bottom in the figure) with a caption head. After obtaining the fused visual tokens $V$ and decoder output query embeddings $Q$ of each scene, the caption head uses $Q$ as caption prefix to identify the described region, and contextual features $V$ surrounding the vote query to complement with more surrounding information for more descriptive caption generation. Finally, the referring/detection boxes are selected from the candidate boxes via the referring scores.
@@ 3 Method
The key idea of 3DGCTR is to harness the relation-oriented capability of the 3DVG task and the object-oriented ability of the 3DDC task to complement each other. The architecture of the adopted 3DVG model is illustrated in Sec. 3.1. We also enhance the feature extraction ability of EDA [34] as our 3DVG model, termed EDA-PMB, which is introduced in Sec. 3.2. To develop the 3DDC capacity, a caption head is appended to the 3DVG model for caption generation, which is detailed in Sec. 3.3. To naturally unify 3DDC and 3DVG tasks in DETR-like architecture, we propose a text prompt for 3DDC as an input to the 3DVG backbone, as described in Sec. 3.4. Last, the training scheme and the inference process are illustrated in Sec. 3.5 and Sec. 3.6, respectively.
@@ 3.1 Preliminaries: 3DVG Model
To ensure the end-to-end training characteristics of our model, we introduce EDA [34] with the setting of single-stage as the basic 3DVG model. EDA is a typical DETR-like [4] model composed of a backbone, a cross-encoder, and an object decoder.
Visual and Textual Backbone. As shown in Figure. 2, PointMetaBased [21] produces the visual tokens $V_{0}\in\mathbb{R}^{n\times d}$ from the input point cloud $P\in\mathbb{R}^{N\times 3}$, while Roberta-base [22] processes referential text for text tokens $T_{0}\in\mathbb{R}^{l\times d}$, where $n$ denotes the number of visual token, $l$ is the number of text token, $N$ denotes the point cloud size, and $d$ is the feature dimension.
Cross-Encoder. The visual tokens $V_{0}$ and text tokens $T_{0}$ enter a dual-pathway cross-encoder alternatively stacking with self-attention, cross-attention, and FFN layers, producing thoroughly fused features $V\in\mathbb{R}^{n\times d}$ and $T\in\mathbb{R}^{l\times d}$. The output visual tokens $V\in\mathbb{R}^{n\times d}$ is input to the KPS module proposed by [23]. The top-k object queries $Q_{0}\in\mathbb{R}^{k\times d}$ are chosen and then fed into the object decoder, caption head and the query scoring branch.
@@ 3.2 Integrating PointMetaBase into EDA
A more powerful visual backbone can better extract the attributes and relative position relationships of objects in 3D scenes, which is helpful for more fine-grained interaction with language. Hence, we replaced the PointNet++ [26] in EDA [34] with PointMetaBase [21]. The main contribution of PointMetaBase is to abstract the point cloud feature extraction network into a Meta Architecture consisting of four key modules and explore the optimal design of each module. However, as the K-Nearest query is used in the module Neighbor Update, the number of downsampled point clouds output by the backbone is not fixed, and thus can’t be used as the visual tokens $V_{0}$ directly. To solve this problem, we design a parameter-free visual token query module. We first use the farthest point sampling (FPS) algorithm to sample a fixed number of $n$ points from the original point cloud $P$ as the candidates $C$:
@@
Then, given the output tokens $V^{\prime}\in\mathbb{R}^{n^{\prime}\times 3}$ from the PointMetaBase, we employ the ball query [26] with a ball radius of $r$ to find $k_{q}$ nearest neighbor tokens in the coordinate space of $V^{\prime}$ for each candidate, then assign the features of $V^{\prime}$ to them using max-pooling. Finally, we can obtain the visual tokens $V_{0}\in\mathbb{R}^{n\times 3}$:
@@
3.3 Caption Head
We argue that the key to unambiguous detailed caption generation is to obtain the relationship between the target object and its close surroundings. As described in Vote2Cap-DETR [8], the vote queries fail to provide adequate attribute and spatial relations. To address this issue, they proposed a caption head named Dual-Clued Captioner (DCC), which introduced the vote queries’ $k$ nearest local context token features as their local surroundings and keys for cross attention. In our method, the top-k object queries $Q_{0}$ generated in the KPS module also have the same problem as the vote queries generated in Vote2Cap-DETR. However, different from Vote2Cap-DETR, where vote queries are constructed based on spatial bias and content information for better caption generation, our KPS-based method focuses on the precise localization of key points of objects to tackle both 3DVG and 3DDC tasks simultaneously. Therefore, we use DCC as our caption head with some modifications to fit this difference. Instead of using vote queries’ $k$ nearest for cross attention, we introduce the visual tokens $V$ that integrate the broader scene context as the object’s interaction and relation information with its surroundings, leading to more contextually nuanced captions.
As shown in Figure 2, DCC is a lightweight transformer decoder-based model comprising two identical transformer decoder blocks, sinusoidal position embeddings $PE(\cdot)$, and a linear classification head. For effective caption generation, DCC processes dual visual cues $V_{c}=(Q,V)$. Firstly, following [32], in captioning a proposal, the standard ’Start Of Sequences’ (SOS) prefix is replaced with the query $Q$ from the described query, identifying the object in focus. Then, for each query embedding $q\in Q$, DDC gives its corresponding dense caption. The masked self-attention and cross-attention mechanism of the transformer decodes the contextual relationship between the query and the visual scene, which can be formulated as: